Explain how the energy, frequency, and wavelength of radiation are related. Give appropriate formula(s).
Answer #1

Let energy of Radiation is E, frequency is f and wavelength is L.

Then, E = (h * c) / L and E = h * f

Where c is the velocity of light.

And c = f * L

These are the 3 formulas.